OpenStudy (anonymous):

Solve the following system of equations: 2x - 5y =22 -2x - 3y = 6 Answer a. (1, -8) b. (-9, -8) c. (-9, 3) d. (8, 9)

jimthompson5910 (jim_thompson5910):

Add the two equations to get 2x - 5y =22 -2x - 3y = 6 ----------- 0x - 8y = 28 So -8y = 28 Solve this for y. Then use this solution to find x.

OpenStudy (anonymous):

@jim_thompson5910 that value isn't in the answer set though. :) I think something may be missing from the equations.

jimthompson5910 (jim_thompson5910):

hmm that's odd...

jimthompson5910 (jim_thompson5910):

probably a typo in the original equations?